# Fourier Circuits in Neural Networks: Unlocking the Potential of Large Language Models in Mathematical Reasoning and Modular Arithmetic

### v2

@@ 4.1 Technique Overview
In this section, we propose techniques overview of the proof for our main result. We use $\mathbf{i}$ to denote $\sqrt{-1}$. Let $f:\mathbb{Z}_{p}\rightarrow\mathbb{C}$. Then, for each frequency $j\in\mathbb{Z}_{p}$, we define $f$ discrete Fourier transform (DFT) as $\widehat{f}(j):=\sum_{\zeta\in\mathbb{Z}_{p}}f(\zeta)\exp(-2\pi\mathbf{i}\cdot j\zeta/p).$ Let $\Omega_{q}^{{}^{\prime}*}$ be the single neuron class-weighted maximum-margin solution set (formally defined in Definition E.6). We first show how we get $\Omega^{\prime*}_{q}$.
Lemma 4.2 (Informal version of Lemma E.8).
If for any $\zeta\in\{1,\ldots,\frac{p-1}{2}\}$, there exists a scaling constant $\beta\in\mathbb{R}$, such that $u_{i}(a_{i})=\beta\cdot\cos(\theta_{u_{i}}^{*}+2\pi\zeta a_{i}/p)$ for any $i\in[k]$ and $w(c)=\beta\cdot\cos(\theta_{w}^{*}+2\pi\zeta c/p)$, where $\theta_{u_{1}}^{*},\dots,\theta_{u_{k}}^{*},\theta_{w}^{*}\in\mathbb{R}$ are some phase offsets satisfying $\theta_{u_{1}}^{*}+\dots+\theta_{u_{k}}^{*}=\theta_{w}^{*}$. Then, we have the following
$\Omega_{q}^{\prime*}=$$\{(u_{1},\dots,u_{k},w)\},\quad\gamma^{*}=\frac{2(k!)}{(2k+2)^{(k+1)/2}(p-1)p^{(k-1)/2}}.$
@@
See formal proof in Appendix E.5. The proof establishes the maximum-margin solution’s sparsity in the Fourier domain through several key steps. Initially, by Lemma E.7, focus is directed to maximizing Eq. (E.4). For odd $p$, Eq. (E.4) can be reformulated with magnitudes and phases of $\widehat{u}_{i}$ and $\widehat{w}$ (discrete Fourie transform of $u_{i}$ and $w$), leading to an equation involving cosine of their phase differences. Plancherel’s theorem is then employed to translate the norm constraint to the Fourier domain. This allows for the optimization of the cosine term in the sum, effectively reducing the problem to maximizing the product of magnitudes of $\widehat{u}_{i}$ and $\widehat{w}$ (Eq. (20)). By applying the inequality of arithmetic and geometric means we have an upper bound for the optimization problem. To achieve the upper bound, equal magnitudes are required for all $\widehat{u}_{i}$ and $\widehat{w}$ at a single frequency, leading to Eq. (24). The neurons are finally expressed in the time domain, demonstrating that they assume a specific cosine form with phase offsets satisfying certain conditions. ∎
Next, we show the number of neurons required to solve the problem and the property of these neurons. We demonstrate how to use these neurons to construct the network $\theta^{*}$.
Lemma 4.3 (Informal version of Lemma F.3).
Let $\cos_{\zeta}(x)$ denote $\cos(2\pi\zeta x/p)$. Then, we have the maximum $L_{2,k+1}$-margin solution $\theta^{*}$ will consist of $2^{2k-1}\cdot{p-1\over 2}$ neurons $\theta^{*}_{i}\in\Omega_{q}^{{}^{\prime}*}$ to simulate $\frac{p-1}{2}$ type of cosine computation, each cosine computation is uniquely determined a $\zeta\in\{1,\ldots,\frac{p-1}{2}\}$. In particular, for each $\zeta$ the cosine computation is $\cos_{\zeta}(a_{1}+\dots+a_{k}-c),\forall a_{1},\dots,a_{k},c\in\mathbb{Z}_{p}$.
@@
See formal proof in Appendix F.3. Our goal is to show that $2^{2k-1}\cdot{p-1\over 2}$ neurons $\theta_{i}^{*}\in\Omega_{q}^{{}^{\prime}*}$ are able to simulate $\frac{p-1}{2}$ type of $\cos$ computation.
We have the following expansion function of $\cos_{\zeta}(x)$, which denotes $\cos(2\pi\zeta x/p)$.
$\cos_{\zeta}(\sum_{i=1}^{k}a_{i})=$ $~{}\sum_{b\in\{0,1\}^{k}}\prod_{i=1}^{k}\cos^{1-b_{i}}(a_{i})\cdot\sin^{b_{i}}(a_{i})\cdot{\bf 1}[\sum_{i=1}^{k}b_{i}\%2=0]\cdot(-1)^{{\bf 1}[\sum_{i=1}^{k}b_{i}\%4=2]}.$
The above equation can decompose a $\cos(\sum)$ to some basic elements. Note that we have $2^{k}$ terms in the above equation. By using the following fact in Lemma F.1,
@@
where each term can be constructed by $2^{k-1}$ neurons. Therefore, we need $2^{k-1}2^{k}$ total neurons. To simulate $\frac{p-1}{2}$ type of simulation, we need $2^{2k-1}\frac{p-1}{2}$ neurons. Then, using the Lemma D.1, we construct the network $\theta^{*}$. Finally, by using the Lemma D.2 from [64], we know that it is the maximum-margin solution. ∎
Now, we are ready to prove our main results.
Proof sketch of Theorem 4.1.
See formal proof in Appendix H.2. By Lemma 4.2, we get $\gamma^{*}$ and the single-neuron class-weighted maximum-margin solution set $\Omega_{q}^{{}^{\prime}*}$. By satisfying Condition 3.8, we know it is used in the maximum-margin solution. By Lemma 4.3, we can construct the network $\theta^{*}$ that uses neurons in $\Omega_{q}^{{}^{\prime}*}$. By Lemma D.2, we know that it is the maximum-margin solution. Finally, by Lemma G.2, we know that all frequencies are covered. ∎
@@ 5 Experiments
In Section 5.1, we conduct simulation experiments to verify our analysis for $k=3,4,5$. In Section 5.2, we show that the one-layer transformer learns 2-dimensional cosine functions in their attention weights. In Section 5.3, we show the grokking phenomenon under different $k$. Please refer to Appendix I.1 for more details about implementation.
Figure 2: All Fourier spectrum frequencies being covered and the maximum normalized power of the embeddings (hidden layer weights). The one-hidden layer network with $m=2944$ neurons is trained on $k=4$-sum mod-$p=47$ addition dataset. We denote $\widehat{u}[i]$ as the Fourier transform of $u[i]$. Let $\max_{i}|\widehat{u}[i]|^{2}/(\sum|\widehat{u}[j]|^{2})$ be the maximum normalized power. Mapping each neuron to its maximum normalized power frequency, (a) shows the final frequency distribution of the embeddings. Similar to our construction analysis in Lemma 4.3, we have an almost uniform distribution over all frequencies. (b) shows the maximum normalized power of the neural network with random initialization. (c) shows, in frequency space, the embeddings of the final trained network are one-sparse, i.e., maximum normalized power being almost 1 for all neurons. This is consistent with our max-margin analysis results in Lemma 4.3. See Figure 6 and Figure 8 in Appendix I.2 for similar results when $k$ is 3 or 5.
5.1 One-hidden Layer Neural Network
@@
We conduct simulation experiments to verify our analysis. In Figure 1 and Figure 2, we use SGD to train a two-layer network with $m=2944=2^{2k-2}\cdot(p-1)$ neurons, i.e., Eq. (2), on $k=4$-sum mod-$p=47$ addition dataset, i.e., Eq. (1). Figure 1 shows that the networks trained with SGD have single-frequency hidden neurons, which support our analysis in Lemma 4.2. Furthermore, Figure 2 demonstrates that the network will learn all frequencies in the Fourier spectrum which is consistent with our analysis in Lemma 4.3. Together, they verify our main results in Theorem 4.1 and show that the network trained by SGD prefers to learn Fourier-based circuits. There are more similar results when $k$ is 3 or 5 in Appendix I.2.
5.2 One-layer Transformer
We find similar results in the one-layer transformer. Recall that the $m$-heads attention layer can be written as
$W^{P}\begin{pmatrix}W_{1}^{V\top}E\cdot\operatorname{softmax}\left({E^{\top}W_{1}^{K}W_{1}^{Q\top}E}\right)\\ \dots\\ W_{m}^{V\top}E\cdot\operatorname{softmax}\left({E^{\top}W_{m}^{K}W_{m}^{Q\top}E}\right)\end{pmatrix},$
where $E$ is input embedding and $W^{P},W^{V},W^{K},W^{Q}$ are projection, value, key and query matrix. We denote $W^{K}W^{Q\top}$ as $W^{KQ}$ and call it attention matrix.
In Figure 3 , we train a one-layer transformer with $m=160$ heads attention, i.e., above equation, on $k=4$-sum mod-$p=31$ addition dataset, i.e., Eq. (1). Figure 3 shows that the SGD trained one-layer transformer learns 2-dim cosine shape attention matrices, which is similar to the one-hidden layer neural networks in Figure 1. This means that the attention layer has a learning mechanism similar to neural networks in the modular arithmetic task. It prefers to learn (2-dim) Fourier-based circuits when trained by SGD. There are more similar results when $k$ is 3 or 5 in Appendix I.3.
@@ 5.2 One-layer Transformer
Figure 3: 2-dimension cosine shape of the trained $W^{KQ}$ (attention weights) and their Fourier power spectrum. The one-layer transformer with attention heads $m=160$ is trained on $k=4$-sum mod-$p=31$ addition dataset. We even split the whole datasets ($p^{k}=31^{4}$ data points) into training and test datasets. Every row represents a random attention head from the transformer. The left figure shows the final trained attention weights being an apparent 2-dim cosine shape. The right figure shows their 2-dim Fourier power spectrum. The results in the figures are consistent with Figure 1. See Figure 9 and Figure 10 in Appendix I.3 for similar results when $k$ is 3 or 5.
5.3 Grokking under Different $k$
Following the experiments’ protocol in [77], we show there is the grokking phenomenon under different $k$. We train two-layer transformers with $m=160$ attention heads on $k=2,3,4,5$-sum mod-$p=97,31,11,5$ addition dataset with 50% of the data in the training. We use different $p$ to guarantee the dataset sizes are roughly equal to each other.
Figure 4 shows that the grokking weakens as the number of $k$ increases, which is consistent with our analysis. It implies that when the ground-truth function class becomes “complicated”, the transformers need to train more steps to fit the training datasets and the generalization tends to be better. Brilliant recent works by [55, 52] argue that, during learning, the network will be first in the lazy training/NTK regime and then transfer to the rich/feature learning regime sharply, leading to a grokking phenomenon. We use the learning steps required for regime switch as a metric of grokking strength.
“Underfitting” in NTK but “overfitting” in feature learning.
Note that NTK is a notorious overparameterized regime, which probably needs a much larger number of neurons than our max-margin convergence case, i.e., much larger than $\Omega(2^{2k})$ in Theorem 4.1. Thus, under the fixed $m$ and increasing $k$, the model may easily escape the NTK regime, or there is no longer an NTK regime. Thus, we will see a weaker grokking phenomenon as the learning steps needed to transfer from the NTK regime to the feature learning regime become fewer. With increasing $k$, the model will have an “underfitting” issue in the NTK regime, meaning the model must need feature learning to fit the task but cannot only fit the task by NTK. However, the model still has an “overfitting” in the feature learning regime.
@@ 5.3 Grokking under Different $k$
Figure 4: Grokking (models abruptly transition from bad generalization to perfect generalization after a large number of training steps) under learning modular addition involving $k=2,3,4,5$ inputs. We train two-layer transformers with $m=160$ attention heads on $k=2,3,4,5$-sum mod-$p=97,31,11,5$ addition dataset with $50$% of the data in the training set under AdamW [54] optimizer 1e-3 learning rate and 1e-3 weight decay. We use different $p$ to guarantee the dataset sizes are roughly equal to each other. The blue curves show training accuracy and the red ones show validation accuracy. There is a grokking phenomenon in all figures. However, as $k$ increases, the grokking phenomenon becomes weak. See explanation in Section 5.3.
@@ 6 Discussion
The interpretability of grokking in large language models (LLMs) is explored in [70]. By examining various intermediate states within the residual stream of the Transformer model, it is validated that the model employs Fourier features to tackle the modular addition task. However, fully comprehending how the Transformer model and LLMs perform modular addition remains challenging based on the current work, particularly from a theoretical standpoint. We contend that beginning with a simplistic model setup and achieving a thorough and theoretical understanding of how the network utilizes Fourier features to address the problem serves as a valuable starting point and it provides a theoretical understanding of the grokking phenomenon. We believe that further study on LLMs will be an interesting and important future direction.
Grokking, Benign Overfitting, and Implicit Bias.
@@
Recently, [115] connects the grokking phenomenon to benign overfitting [9, 13, 99, 26, 28]. It shows how the network undergoes a grokking period from catastrophic to benign overfitting. [55, 52] uses implicit bias [85, 32, 48, 92, 69, 12, 59, 44, 113, 114] to explain grokking, where grokking happens if the early phase bias implies an overfitting solution while late phase bias implies a generalizable solution. The intuition from the benign overfitting and the implicit bias well align with our observation in Section 5.3. It is interesting and valuable to rigorously analyze the grokking or emergent ability under different function class complexities, e.g., Eq (1). We leave this challenge problem as a future work.
Connection to Parity and SQ Hardness.
@@
If we let $p=2$, then $(a_{1}+\dots+a_{k})\bmod p$ will degenerate to parity function, i.e., $b_{1},\dots,b_{k}\in\{\pm 1\}$ and determining $\prod_{i=1}^{k}b_{i}$. Parity functions serve as a fundamental set of learning challenges in computational learning theory, often used to demonstrate computational obstacles [89]. In particular, $(n,k)$-sparse parity problem is notorious hard to learn, i.e., Statistical Query (SQ) hardness [8]. [23] showed that one-hidden layer networks need an $\Omega(\exp(k))$ number of neurons or an $\Omega(\exp(k))$ number of training steps to successfully learn it by SGD. In our work, we are studying Eq. (2), which is a more general function than parity and indeed is a learning hardness. Our Theorem 4.1 states that we need $\Omega(\exp(k))$ number of neurons to represent the maximum-margin solution, which well aligns with existing works. Our experiential results in Section 5.1 are also consistent. Hence, our modular addition involving $k$ inputs function class is a good data model to analyze and test the model learning ability, i.e., approximation power, optimization, and generalization.
High Order Correlation Attention.
@@
[86, 3, 4] state that, when $k=3$, $a_{1}+a_{2}+a_{3}\bmod p$ is hard to be captured by traditional attention. Thus, they introduce high-order attention to capture high-order correlation from the input sequence. However, in Section 5.2, we show that one-layer transformers have a strong learning ability and can successfully learn modular arithmetic tasks even when $k=5$. This implies that the traditional attention may be more powerful than we expect. We conjecture that the layer norm and residual connection contribute as they are ignored by most transformer learning theoretical analysis work [46, 62, 58, 105].

### v3

@@ 4.1 TECHNIQUE OVERVIEW
In this section, we propose techniques overview of the proof for our main result. We use $\mathbf{i}$ to denote $\sqrt{-1}$. Let $f:\mathbb{Z}_{p}\rightarrow\mathbb{C}$. Then, for each frequency $j\in\mathbb{Z}_{p}$, we define $f$ discrete Fourier transform (DFT) as $\widehat{f}(j):=\sum_{\zeta\in\mathbb{Z}_{p}}f(\zeta)\exp(-2\pi\mathbf{i}\cdot j\zeta/p).$ Let $\Omega_{q}^{{}^{\prime}*}$ be the single neuron class-weighted maximum-margin solution set (formally defined in Definition F.6). We first show how we get $\Omega^{\prime*}_{q}$.
Lemma 4.2 (Informal version of Lemma F.8).
If for any $\zeta\in\{1,\ldots,\frac{p-1}{2}\}$, there exists a scaling constant $\beta\in\mathbb{R}$, such that $u_{i}(a_{i})=\beta\cdot\cos(\theta_{u_{i}}^{*}+2\pi\zeta a_{i}/p)$ for any $i\in[k]$ and $w(c)=\beta\cdot\cos(\theta_{w}^{*}+2\pi\zeta c/p)$, where $\theta_{u_{1}}^{*},\dots,\theta_{u_{k}}^{*},\theta_{w}^{*}\in\mathbb{R}$ are some phase offsets satisfying $\theta_{u_{1}}^{*}+\dots+\theta_{u_{k}}^{*}=\theta_{w}^{*}$. Then, we have $\Omega_{q}^{\prime*}=\{(u_{1},\dots,u_{k},w)\},$ and $\gamma^{*}=\frac{2(k!)}{(2k+2)^{(k+1)/2}(p-1)p^{(k-1)/2}}$.
@@
See formal proof in Appendix F.5. The proof establishes the maximum-margin solution’s sparsity in the Fourier domain through several key steps. Initially, by Lemma F.7, focus is directed to maximizing Eq. (F.4). For odd $p$, Eq. (F.4) can be reformulated with magnitudes and phases of $\widehat{u}_{i}$ and $\widehat{w}$ (discrete Fourier transform of $u_{i}$ and $w$), leading to an equation involving cosine of their phase differences. Plancherel’s theorem is then employed to translate the norm constraint to the Fourier domain. This allows for the optimization of the cosine term in the sum, effectively reducing the problem to maximizing the product of magnitudes of $\widehat{u}_{i}$ and $\widehat{w}$ (Eq. (20)). By applying the inequality of arithmetic and geometric means, we have an upper bound for the optimization problem. To achieve the upper bound, equal magnitudes are required for all $\widehat{u}_{i}$ and $\widehat{w}$ at a single frequency, leading to Eq. (24). The neurons are finally expressed in the time domain, demonstrating that they assume a specific cosine form with phase offsets satisfying certain conditions. ∎
Next, we show the number of neurons required to solve the problem and the properties of these neurons. We demonstrate how to use these neurons to construct the network $\theta^{*}$.
Lemma 4.3 (Informal version of Lemma G.3).
Let $\cos_{\zeta}(x)$ denote $\cos(2\pi\zeta x/p)$. Then, we have the maximum $L_{2,k+1}$-margin solution $\theta^{*}$ will consist of $2^{2k-1}\cdot{p-1\over 2}$ neurons $\theta^{*}_{i}\in\Omega_{q}^{{}^{\prime}*}$ to simulate $\frac{p-1}{2}$ type of cosine computation, where each cosine computation is uniquely determined a $\zeta\in\{1,\ldots,\frac{p-1}{2}\}$. In particular, for each $\zeta$ the cosine computation is $\cos_{\zeta}(a_{1}+\dots+a_{k}-c),\forall a_{1},\dots,a_{k},c\in\mathbb{Z}_{p}$.
@@
See formal proof in Appendix G.3. Our goal is to show that $2^{2k-1}\cdot{p-1\over 2}$ neurons $\theta_{i}^{*}\in\Omega_{q}^{{}^{\prime}*}$ are able to simulate $\frac{p-1}{2}$ type of $\cos$ computation. We have the following expansion function of $\cos_{\zeta}(x)$, which denotes $\cos(2\pi\zeta x/p)$.
$\cos_{\zeta}(\sum_{i=1}^{k}a_{i})=$ $~{}\sum_{b\in\{0,1\}^{k}}\prod_{i=1}^{k}\cos^{1-b_{i}}(a_{i})\cdot\sin^{b_{i}}(a_{i})$
$~{}\cdot{\bf 1}[\sum_{i=1}^{k}b_{i}\%2=0]\cdot(-1)^{{\bf 1}[\sum_{i=1}^{k}b_{i}\%4=2]}.$
The above equation can decompose a $\cos(\sum)$ to some basic elements. We have $2^{k}$ terms in the above equation. By using the following fact in Lemma G.1,
@@
where each term can be constructed by $2^{k-1}$ neurons. Therefore, we need $2^{k-1}2^{k}$ total neurons. To simulate $\frac{p-1}{2}$ type of simulation, we need $2^{2k-1}\frac{p-1}{2}$ neurons. Then, using the Lemma E.1, we construct the network $\theta^{*}$. By using the Lemma E.2 from [58], we get it is the maximum-margin solution. ∎
@@ 5 EXPERIMENTS
First, we conduct simulation experiments to verify our analysis for $k=3,4,5$. Then, we show that the one-layer transformer learns 2-dimensional cosine functions in their attention weights. Finally, we show the grokking phenomenon under different $k$. Please refer to Appendix J.1 for details about implementation.
@@
One-hidden Layer Neural Network. We conduct simulation experiments to verify our analysis. In Figure 1 and Figure 2, we use SGD to train a two-layer network with $m=2944=2^{2k-2}\cdot(p-1)$ neurons, i.e., Eq. (2), on $k=4$-sum mod-$p=47$ addition dataset, i.e., Eq. (1). Figure 1 shows that the networks trained with SGD have single-frequency hidden neurons, which support our analysis in Lemma 4.2. Furthermore, Figure 2 demonstrates that the network will learn all frequencies in the Fourier spectrum which is consistent with our analysis in Lemma 4.3. Together, they verify our main results in Theorem 4.1 and show that the network trained by SGD prefers to learn Fourier-based circuits. There are more similar results when $k$ is 3 or 5 in Appendix J.2.
@@
Figure 3: 2-dimension cosine shape of the trained $W^{KQ}$ (attention weights) and their Fourier power spectrum. The one-layer transformer with attention heads $m=160$ is trained on $k=4$-sum mod-$p=31$ addition dataset. We even split the whole datasets ($p^{k}=31^{4}$ data points) into training and test datasets. Every row represents a random attention head from the transformer. The left figure shows the final trained attention weights being an apparent 2-dim cosine shape. The right figure shows their 2-dim Fourier power spectrum. The results in the figures are consistent with Figure 1. See Figure 9 and Figure 10 in Appendix J.3 for similar results when $k$ is 3 or 5.
One-layer Transformer. We find similar results in one-layer transformers. Let $E$ be input embedding and $W^{P},W^{V},W^{K},W^{Q}$ be projection, value, key and query matrix. The $m$-heads attention layer can be written as
$W^{P}\begin{pmatrix}W_{1}^{V\top}E\cdot\operatorname{softmax}\left({E^{\top}W_{1}^{K}W_{1}^{Q\top}E}\right)\\ \dots\\ W_{m}^{V\top}E\cdot\operatorname{softmax}\left({E^{\top}W_{m}^{K}W_{m}^{Q\top}E}\right)\end{pmatrix}.$
We denote $W^{K}W^{Q\top}$ as $W^{KQ}$ and call it attention matrix. In Figure 3 , we train a one-layer transformer with $m=160$ heads attention and hidden dimension 128, i.e., above equation, on $k=4$-sum mod-$p=31$ addition dataset, i.e., Eq. (1). Figure 3 shows that the SGD trained one-layer transformer learns 2-dim cosine shape attention matrices, which is similar to the one-hidden layer neural networks in Figure 1. This means that the attention layer has a learning mechanism similar to neural networks in the modular arithmetic task. It prefers to learn (2-dim) Fourier-based circuits when trained by SGD. There are more similar results when $k$ is 3 or 5 in Appendix J.3.
@@
Figure 4: Grokking (models abruptly transition from bad generalization to perfect generalization after a large number of training steps) under learning modular addition involving $k=2,3,4,5$ inputs. We train two-layer transformers with $m=160$ attention heads on $k=2,3,4,5$-sum mod-$p=97,31,11,5$ addition dataset with $50$% of the data in the training set under AdamW [48] optimizer 1e-3 learning rate and 1e-3 weight decay. We use different $p$ to guarantee the dataset sizes are roughly equal to each other. The blue curves show training accuracy and the red ones show validation accuracy. There is a grokking phenomenon in all figures. However, as $k$ increases, the grokking phenomenon becomes weak. See explanation in Section 5.
Grokking under Different $k$. To support our data setting importance, we study grokking phenomenon on our data distribution. Following the experiments’ protocol in [69], we show there is the grokking phenomenon under different $k$. We train two-layer transformers with $m=160$ attention heads and hidden dimension as 128 on $k=2,3,4,5$-sum mod-$p=97,31,11,5$ addition dataset with 50% of the data in training. We use different $p$ to guarantee the dataset sizes are roughly equal to each other. Figure 4 shows that the grokking weakens as the number of $k$ increases, which is consistent with our analysis. When $k$ increases, the function class will become more complicated, as we may need more neurons to achieve the max-margin solution. Thus, we use our Theorem 4.1 as a metric to measure the data complexity. It implies that when the ground-truth function class becomes “complicated”, the transformers need to train more steps to fit the training datasets, and the generalization tends to be better. Brilliant recent works by [49, 45] argue that, during learning, the network will be first in the lazy training/NTK regime and then transfer to the rich/feature learning regime sharply, leading to a grokking phenomenon. We use learning steps required for regime switch as a metric of grokking strength.
NTK is a notorious overparameterized regime, which probably needs a much larger number of neurons than our max-margin convergence case, i.e., much larger than $\Omega(2^{2k})$ in Theorem 4.1. Thus, under the fixed $m$ and increasing $k$, the model may easily escape the NTK regime, or there is no longer an NTK regime. Thus, we will see a weaker grokking phenomenon as the learning steps needed to transfer from the NTK regime to the feature learning regime become fewer. With increasing $k$, the model will have an “underfitting” issue in the NTK regime, meaning the model must need feature learning to fit the task but cannot only fit the task by NTK. However, the model still has an “overfitting” in the feature learning regime.
@@ 6 DISCUSSION
Grokking in Transformers. The interpretability of grokking in Transformers is explored in [63]. By examining various intermediate states within the residual stream of the Transformer model, it is validated that the model employs Fourier features to tackle the modular addition task. However, fully comprehending how the Transformer model and LLMs perform modular addition remains challenging based on the current work, particularly from a theoretical standpoint. We contend that beginning with a simplistic model setup and achieving a thorough and theoretical understanding of how the network utilizes Fourier features to address the problem serves as a valuable starting point and it provides a theoretical understanding of the grokking phenomenon. We believe that further study on Transformers will be an interesting and important future direction.
@@
Grokking, Benign Overfitting, and Implicit Bias. Recently, [103] connects the grokking phenomenon to benign overfitting [9, 13, 90, 25, 27]. It shows how the network undergoes a grokking period from catastrophic to benign overfitting. [49, 45] uses implicit bias [75, 29, 41, 83, 62, 12, 53, 38, 101, 102] to explain grokking, where grokking happens if the early phase bias implies an overfitting solution while late phase bias implies a generalizable solution. The intuition from the benign overfitting and the implicit bias well align with our observation in Section 5. It is interesting and valuable to rigorously analyze the grokking or emergent ability under different function class complexities, e.g., Eq (1). We leave this challenge problem as a future work.
@@
Connection to Parity and SQ Hardness. If we let $p=2$, then $(a_{1}+\dots+a_{k})\bmod p$ will degenerate to parity function, i.e., $b_{1},\dots,b_{k}\in\{\pm 1\}$ and determining $\prod_{i=1}^{k}b_{i}$. Parity functions serve as a fundamental set of learning challenges in computational learning theory, often used to demonstrate computational obstacles [80]. In particular, $(n,k)$-sparse parity problem is notorious hard to learn, i.e., Statistical Query (SQ) hardness [8]. [22] showed that one-hidden layer networks need an $\Omega(\exp(k))$ number of neurons or an $\Omega(\exp(k))$ number of training steps to successfully learn it by SGD. In our work, we are studying Eq. (2), which is a more general function than parity and indeed is a learning hardness. Our Theorem 4.1 states that we need $\Omega(\exp(k))$ number of neurons to represent the maximum-margin solution, which well aligns with existing works. Our experiential results in Section 5 are also consistent. Hence, our modular addition involving $k$ inputs function class is a good data model to analyze and test the model learning ability, i.e., approximation, optimization, and generalization.
@@
High Order Correlation Attention. [76, 3, 4] state that, when $k=3$, $a_{1}+a_{2}+a_{3}\bmod p$ is hard to be captured by traditional attention. Thus, they introduce high-order attention to capture high-order correlation from the input sequence. However, in Section 5, we show that one-layer transformers have a strong learning ability and can successfully learn modular arithmetic tasks even when $k=5$. This implies that the traditional attention may be more powerful than we expect.